OpenStudy (anonymous):

A certain radioactive element has a half-life of one hour. If you start with 1.0 gram of the element at noon, how much of the radioactive element will be left at 2 p.m.? ( 3 points) Answer 0.50 grams 0.25 grams 0.125 grams 0.0 grams

OpenStudy (anonymous):

Nt = No (1/2) ^ (time/halftime) Nt = 1 (1/2)^(2/1) = 1/4 grams